hello tutor, please help me solve this word problem. Thank You.
Mark is now 2 times as old as his son David. 16 years ago, Mark was 6 years older than 3 times as David's age. How old is David now?
Let David be N years old now.
Thank you again! :)

Answer by sachi(548)

Mark is now 2 times as old as his son David.
Let David be N years old now.
then Mark is 2N yrs now
16 years ago,
David was N-16 years old
& Mark was 2N-16 yrs
Mark was 6 years older than 3 times as David's age.
So 2N-16=6+3(N-16)=6+3N-48=3N-42
or 3N-2N=42-16=26
or N= 26 yrs=David's age now
Ans